active ingredient of a pill comprises 45 % of the pill mass. if a pill contains 325 mg of active ingredient, what is the total mass of pill
Answer #1

The active ingredient of a pill comprises 45 % of the pill mass. if a pill contains 325 mg of the active ingredient, what is the total mass of pill

45/100 * x = 325

= 45x = 325 *100

45x = 32500

X = 32500/45

722.222

722.2 mg